OpenStudy (anonymous):

what is P(greater Than 4) ? 1/3 1/2

OpenStudy (anonymous):



OpenStudy (anonymous):

@satellite73 Can You Help Me Please

OpenStudy (anonymous):

If this is a die then each side has an equally likely chance. P(x>4) is therefore P(5) = 1/6 and P(6) = 1/6 so P(x >4 ) = 2/6 = 1/3

OpenStudy (anonymous):

P(x>4) is therefore P(5) = 1/6 and P(6) = 1/6 so P(x >4 ) = 2/6 = 1/3